### v1

# FedIA: A Plug-and-Play Importance-Aware Gradient Pruning Aggregation Method for Domain-Robust Federated Graph Learning on Node Classification

Abstract
Federated Graph Learning (FGL) under domain skew—as observed on platforms such as Twitch Gamers and multilingual Wikipedia networks—drives client models toward incompatible representations, rendering naïve aggregation both unstable and ineffective. We find that the culprit is not the weighting scheme but the noisy gradient signal: empirical analysis of baseline methods suggests that a vast majority of gradient dimensions can be dominated by domain-specific variance. We therefore shift focus from ”aggregation-first” to a projection-first strategy that denoises client updates before they are combined. The proposed FedIA framework realises this Importance-Aware idea through a two-stage, plug-and-play pipeline: (i) a server-side top-$\rho$ mask keeps only the most informative $\sim 5\%$ of coordinates, and (ii) a lightweight influence-regularised momentum weight suppresses outlier clients. FedIA adds no extra uplink traffic and only negligible server memory, making it readily deployable. On both homogeneous (Twitch Gamers) and heterogeneous (Wikipedia) graphs, it yields smoother, more stable convergence and higher final accuracy than nine strong baselines. A convergence sketch further shows that dynamic projection maintains the optimal $\mathcal{O}\!\bigl{(}\sigma^{2}/\sqrt{T}\bigr{)}$ rate. CODE: https://github.com/itoritsu/FedIA.
1 introduction
Figure 1: Overview of FedIA in a domain-skewed FGL scenario. (a) Domain-skewed clients (e.g., institutions from different sources) produce gradients dominated by domain-specific noise, causing divergent global representations. (b) FedIA applies a two-stage server-side filter: a global top-$\rho$ mask preserves approximately 5% of key coordinates (denoising), followed by influence-regularised momentum weighting (outlier suppression). (c) After projection, training curves stabilize and the global model converges consistently, without extra communication or client-side state overhead.
Federated graph learning (FGL) now powers privacy-critical services such as friend recommendation on live-streaming platforms (e.g., Twitch Gamers) and hyperlink prediction in multilingual knowledge graphs (e.g., Wikipedia Network). Real-world graph data across clients frequently exhibits severe domain skew due to divergent node features and structural patterns arising from geographic or linguistic differences, as illustrated in Figure 1. Such domain discrepancies inevitably cause client gradients to contain substantial domain-specific noise, undermining the consistency of global model representations and severely hindering convergence.
In such domain-skewed scenarios, naïve aggregation schemes rapidly degrade into unstable optimization. Figure 1 visualizes this issue: gradients from domain-skewed clients push the global model parameters toward incompatible regions, resulting in high variance and oscillations during training. As our empirical evidence further shows, standard federated averaging (FedAvg) and even state-of-the-art aggregation approaches remain unstable, unable to achieve reliable and robust convergence.
Prior solutions primarily fall into two impractical categories. Parameter regularization methods (e.g., FedProx (Li et al. 2020), FedDyn (Acar et al. 2021)) attempt to penalize deviations from global parameters but fail to explicitly handle gradient-level noise. Prototype-based or mask-based methods (e.g., FGGP (Wan, Huang, and Ye 2024), FedHEAL (Chen, Huang, and Ye 2024)) attempt to synchronize domain representations by exchanging prototypes or maintaining historical client masks, but incur prohibitive communication and memory overheads, respectively. Critically, none of these methods tackles the root cause—gradient noise—prior to aggregation.
We argue that aggregation is not the primary issue to solve; pre-conditioning the gradient signals is. Our analysis on baseline models like FedAvg reveals that a significant portion of gradient coordinates, potentially over 90% in some cases, can carry domain-specific noise, and retaining a small subset (around 5%) suffices to maintain robust convergence. Moreover, we emphasize that directly transmitting explicit gradients—rather than weight updates—eliminates server-side inference noise, further ensuring high-quality gradient signals for aggregation. This insight motivates a fundamental paradigm shift toward a projection-dominant approach, prioritizing gradient filtering before aggregation.
FedIA instantiates this view through a two-stage, zero-communication-overhead pipeline, as depicted in Figure 1. In Stage 1, FedIA employs a global importance mask to preserve only the top-$\rho$ fraction of gradient coordinates, effectively denoising the aggregated gradient space. In Stage 2, a lightweight influence-regularised momentum weighting further stabilizes optimization by dampening the influence of outlier clients. Importantly, FedIA introduces no additional communication or client-side memory overhead, making it practical and easy to deploy.
This paper makes threefold contributions:
  • Practicality: We propose FedIA, a zero-communication-overhead, drop-in aggregation framework that effectively stabilizes federated graph learning across realistic scenarios including both homogeneous (Twitch Gamers) and heterogeneous (Wikipedia Network) graphs.
  • Framework: We introduce a hierarchical gradient pre-conditioning paradigm, and provide a supporting convergence analysis that connects our method to established theoretical guarantees.
  • Evidence: Comprehensive experiments and qualitative privacy discussions validate the efficacy and practical advantages of FedIA. The source code and experimental setup will be publicly released.
The remainder of this paper is structured as follows. Section 2 formalizes the basic background of FGL and the limitations of previous works; Section 3 details the challenges and core insights; Section 4 describes FedIA in depth; Section 5 presents crucial discussion on communication, deployment, and privacy; Section 6 analyzes the effectiveness and robustness of FedIA; and Section 7 summarizes the contributions of this work. Due to space limits, a further discussion with an empirical study is provided in Appendix A, and more extensive experiments are presented in Appendix B.
2 Background
2.1 Preliminary
In Preliminary, we provide a basic introduction to the research scenario.
Graph Fundamentals. A graph is defined as $\mathcal{G}=(V,E,\mathbf{X})$, where $V={v_{1},...,v_{n}}$ denotes the node set, $E\subseteq V\times V$ represents the edge set, and $\mathbf{X}\in\mathbb{R}^{n\times d_{\text{feat}}}$ is the node feature matrix with feature dimension $d_{\text{feat}}$. For semi-supervised node classification tasks, only a subset of nodes has labels, denoted by $\mathbf{Y}\in\mathbb{R}^{n\times c}$, with $c$ being the number of classes.
Graph Neural Networks (GNNs). GNNs generate node embeddings through iterative message passing, where the embedding $\mathbf{h}_{i}^{(l)}$ for node $v_{i}$ at the $l$-th layer is computed as:
$\mathbf{h}_{i}^{(l)}=\sigma\left(\mathbf{W}^{(l)}\cdot\text{AGGREGATE}\left({\mathbf{h}_{j}^{(l-1)}\mid v_{j}\in\mathcal{N}(v_{i})}\right)\right),$ (1)
where $\mathcal{N}(v_{i})$ denotes the neighbors of node $v_{i}$, $\sigma(\cdot)$ is an activation function, and $\text{AGGREGATE}(\cdot)$ is a differentiable, permutation-invariant function. Final node representations from the last layer are typically utilized for downstream classification tasks.
Federated Learning (FL). In standard federated learning with $K$ clients, the global model parameters $\mathbf{w}t$ at round $t$ are updated by aggregating local client updates:
$\mathbf{w}{t+1}=\mathbf{w}t-\eta\cdot\frac{1}{K}\sum{k=1}^{K}\nabla\mathcal{L}_{k}(\mathbf{w}_{t}),$ (2)
where $\mathcal{L}_{k}$ denotes the local loss function of client $k$, and $\eta$ is the learning rate.
Federated Graph Learning (FGL) for Node Classification. In the context of FGL for node classification, each client $k$ owns a private graph $\mathcal{G}k=(V_{k},E_{k},\mathbf{X}k)$ with adjacency matrix $\mathbf{A}k$ and node features $\mathbf{X}k$. The GNN encoder $f{\text{GNN}}$ generates node embeddings, which are subsequently classified by $g{\text{cls}}$. The global objective is formulated as:
$\sum{k=1}^{K}\frac{|V_{k}|}{|V|}\mathbb{E}{(\mathbf{X}k,\mathbf{A}k,\mathbf{Y}k)}\left[\mathcal{L}_{\text{task}}(g_{\text{cls}}(f_{\text{GNN}}(\mathbf{X}_{k},\mathbf{A}_{k})),\mathbf{Y}k)\right],$ (3)
where $|V|=\sum_{k}|V_{k}|$ and $\mathcal{L}{\text{task}}$ is the node classification loss.
Domain Skew in Node-level Data. Domain skew occurs when joint distributions of adjacency and node features differ across domains, i.e., $P(\mathbf{A},\mathbf{X})\neq P^{\prime}(\mathbf{A},\mathbf{X})$, even though the label distributions remain stable ($P(\mathbf{Y})=P^{\prime}(\mathbf{Y})$). Such domain discrepancies challenge FGL models’ ability to generalize effectively across heterogeneous clients.
2.2 Related Works
In Related Works, we present the shortcomings of existing methods from multiple perspectives.
Federated Graph Learning Approaches. FGL combines FL and GNN to leverage distributed graph data while preserving privacy. Existing methods broadly include FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), FedOPT (Reddi et al. 2020), FedDyn (Acar et al. 2021), and MOON (Li, He, and Song 2021), originally developed for image classification, which directly aggregate parameters or rely on model regularization to enhance stability. However, these parameter-based approaches inherently lack mechanisms to resolve domain-induced discrepancies explicitly, thus exhibiting degraded performance under severe graph domain skew.
Prototype-based Approaches. Recent works such as FGGP (Wan, Huang, and Ye 2024), FedProc (Mu et al. 2023) and FedProto (Tan et al. 2022) adopt prototype-based approaches, extracting and exchanging prototypes representing each client’s feature distribution. Although prototypes effectively capture client-specific domain knowledge, they significantly increase communication overhead by continuously transmitting additional prototype vectors, imposing substantial costs on resource-constrained edge devices.
Fairness-oriented Masking. FedHEAL (Chen, Huang, and Ye 2024) and FairFL (Zhang, Kou, and Wang 2020) propose masking strategies to balance domain contributions by maintaining per-round accumulated masks at the server side. Despite effectively mitigating domain bias, such masking strategies lead to prohibitive server-side memory consumption, scaling poorly to scenarios involving numerous clients or larger models.
Gradient-level Sparsification. Recent advances like Rethink-FL (Duesing and Cimiano 2025) and gradient pruning methods (Xue et al. 2024) directly sparsify gradients to enhance efficiency and reduce privacy leakage. However, these approaches primarily target homogeneous CNN-based classification tasks and do not explicitly address the unique structural and distributional characteristics of graph data in federated scenarios.
These limitations motivate us to explore gradient-level aggregation mechanisms tailored specifically for federated graph learning under node-level domain heterogeneity, addressing both efficiency and domain fairness in a unified manner.
3 Challenges and Insights
Figure 2: Empirical evidence of core challenges in domain-skewed FGL and the effectiveness of our approach. (a) CKA-based visualization reveals severe representation inconsistency between local models and the global model under FedAvg. Our approach (FedAvg+IA) promotes greater uniformity, indicating enhanced domain generalization. (b) Our method achieves a superior final convergence state (lower validation loss) compared to baselines. This is accomplished by operating on a highly sparse gradient subspace, retaining only a core 5.04% of the parameters. (c) The state-of-the-art method, FGGP, demonstrates significant convergence instability across various settings, highlighting the fragility of existing aggregation schemes.
Figure 2 summarises the empirical symptoms we repeatedly observe under domain–skewed federated graph learning (FGL). Two causally–coupled challenges emerge and motivate a shift from classical weight-space aggregation to the gradient-subspace view adopted in this work.
Challenge 1: GNN-specific representation divergence.
When FedAvg aggregates heterogeneous clients, the resulting global encoder aligns almost exclusively with a few dominant domains; the centred-kernel‐alignment (CKA) heat map in Fig. 2(a) reproduces this trend. Graph neural networks exacerbate the problem: (i) message-passing layers broadcast local bias along edges, (ii) batch normalisation amplifies distributional gaps, and (iii) fully-connected heads over-fit class priors of data-rich clients. The global model therefore becomes a patchwork of incompatible sub-representations rather than a truly domain-invariant encoder.
Challenge 2: Variance-dominated gradient noise and rugged optimisation.
Representation divergence manifests as high-variance, often conflicting gradients. Baseline FedAvg oscillates on a rugged loss surface (Fig. 2(b)), while a state-of-the-art aggregator, FGGP, remains unstable across settings (Fig. 2(c)). Simply exchanging larger messages—prototypes, local statistics or second moments—does not tame the noise; the optimiser continues to chase contradictory directions.
Core insight: Projection-dominant aggregation.
Recent analyses of sparsified and projected SGD show that retaining $\leq 10\%$ coordinates can preserve, or even improve, convergence under severe heterogeneity. Our own result in Fig. 2(b) is even more striking: with only 5.04% gradient coordinates retained, the model converges faster and to a markedly lower validation loss. This suggests that most of the high-dimensional gradient space merely carries domain-specific noise. Hence we posit the Projection-dominant Hypothesis: first project every client update onto a global, domain-invariant subspace to suppress variance; only then apply a lightweight, stabilising client-level weighting.
Why explicit gradients?
Uploading weight deltas forces the server to infer the true gradient via $(W^{(t)}-W^{(t+1)})/\eta$ and, for adaptive optimisers, to divide by local second moments—both steps inject estimation noise. By transmitting the explicit gradient, clients expose how their GNN encoders react to domain-specific sub-graph structures, giving the server a high-fidelity signal that can be filtered before aggregation. Section 4 formalises this rationale and turns the above hypothesis into a two-stage algorithm: global top-$\rho$ masking followed by influence-regularised momentum weighting.